OpenStudy (anonymous):

1. 7+3x=5x-9 a.8 b.6 c.-6 2. 7x-8=3x a.-2 b.2 c.10 d.-10 3.4m=-5m-27 a.12 b.27 c.3 d.-3 4. 4(2y-3)=y-8 a.4/7 b.-2/7 c.20/7

OpenStudy (anonymous):

7+3x = 5x-9 7+9 = 5x-3x 16=2x x=8

OpenStudy (anonymous):

try remaining

OpenStudy (anonymous):

whats remaining ?

OpenStudy (anonymous):

2,3 and 4 questions

OpenStudy (anonymous):

Okay i'll try @bhaskarbabu

OpenStudy (anonymous):

2: x=2 3: m=-3 y=4/7